5. Compute the convolution, x[i]*u[n], when x[n] = 2* cos(+7^ )([n – 1] -u[n–6)
Answer #1

xen * ven) uin) = { heidI, I, I, ---} ains cos In) U(n-1) - Vino) 219) = a cosity) = x(2) - 22 Cos (T) = -4 X137 = 23 cos (3- - - - t 19-21)871 +(5-2)87+(227811+ (8-2178h-13:30)8 h =(4) { .--7171 intinaio} = (u7h - -- - 71 71 71 71 71 h h- - - - 0